### v2

@@ Abstract
Local Attention-guided Message Passing Mechanism (LAMP) adopted in Graph Attention Networks (GATs) is designed to adaptively learn the importance of neighboring nodes for better local aggregation on the graph, which is able to effectively bring the representations of similar neighbors closer, thus shows stronger discrimination ability. However, existing GATs suffer from a significant discrimination ability decline in heterophilic graphs because the high proportion of dissimilar neighbors can weaken the self-attention of the central node, jointly resulting in the central node’s deviation from similar nodes in the representation space. This kind of effect generated by neighboring nodes is referred to as Distraction Effect (DE) in this paper. In order to estimate and weaken the DE of neighboring nodes, we propose a Causally graph Attention network for Trimming heterophilic graph (CAT). To estimate the DE, since the DE are generated through two paths (grab the attention assigned to neighbors and reduce the central node’s self-attention), we use Total Effect to model DE, which is a kind of causal estimand and can be estimated from intervened data; To weaken the DE, we identify the neighbors with the highest DE (we call them Distraction Neighbors) and remove them. We adopt three representative GATs as the base model within the proposed CAT framework, and conduct experiments on seven heterophilic datasets in three different sizes. Comparative experiments show that CAT is able to improve the node classification accuracy of all base GAT models. Ablation experiments and visualization further validate the enhancement of discrimination ability brought by CAT. Besides, the framework of CAT is plug-and-play and can be introduced to any LAMP-driven GATs because it learns a trimmed graph in the attention learning stage, instead of modifying model architecture or searching for new neighbors globally. The source code is available at https://github.com/GeoX-Lab/CAT.
@@ 1 Introduction
Graph Neural Network (GNN) is the most reliable and prevailing benchmark model for graph learning. With its effectiveness in representing irregular graph data, GNN achieves state-of-the-art performance in tasks such as node classification, link prediction, graph classification, graph generation and graph similarity calculation [1], and has also been widely applied in recommend system [2, 3, 4, 5, 6], computer vision [7, 8, 9, 10], natural language processing [11], molecular [12, 13, 14, 15], transportation [16, 17, 18, 19], the internet of things [20], epidemiology [21] and other fields. The power of GNN in graph representation primarily stems from its ability to aggregate information, mostly followed the message passing mechanism [22, 23], which can build invariant input representations for the central node based on its local neighbors. Existing GNNs vary in different aggregation operation in accordance with their fundamental assumption about the influence of neighbors, but mostly are built on strong homophily hypothesis of the graph, which obeys the rule that neighbors tend to be similar [24]. Among them Graph Attention Network (GAT) [25] is a representative one, which adaptively learns the importance of neighbors for aggregation through the Local Attention-guided Message Passing Mechanism (LAMP), thereby has the potential for better performance on high homophily graphs. However, the reverse of the coin is a decline when dealing with low homophily graphs. Experiments have shown that GNNs under a strong homophily hypothesis experience significant decline in node classification task [24] when the input graph is heterophilic, and we find that LAMP-driven models like GAT get the most conspicuous decline (as shown in Section 4.1). The main reason is the presence of a high proportion of dissimilar neighbors. Dissimilar neighbors influence the representation of the central node through their assigned attention and weaken the self-attention of the central node, both can result in the central node deviating from its similar nodes in the representation space. We refer to this impact of neighboring nodes on the central node as Distraction Effect (DE), which is generated through two paths (capturing the attention assigned to neighbors and reducing the central node’s self-attention) and leads to the deviation of the representation of central nodes. How to improve the discrimination ability on heterophilic graphs is an important challenge for GATs to overcome.
@@
Existing methods primarily propose solutions for improving the discrimination ability on heterophilic graphs of general GNNs, with only a few offering specific solution for GATs. These approaches can be categorized into two groups: GNN Architecture Tactic and Graph Structure Tactic. GNN Architecture Tactic focuses on modifying the GNN architecture to better utilize the information from neighboring nodes for aggregation. Novel aggregation mechanisms are proposed to adjust the weights of neighbors [26, 27, 28, 29, 30, 31], and some works try to fuse information from different GNN layers in a new way [32, 33, 34, 35]. Self-supervised learning is also adopted to capture more information from neighbors [36, 37, 38]. Graph Structure Tactic, on the other way, focuses on making heterophilic graphs more homophilic by finding more similar nodes for aggregation. These methods tend to look for high-order neighbors [39, 40] or nearest neighbors in some hidden space [41, 42, 43, 44, 45, 46], forcing the central node to be closer to similar nodes in the representation space. Specific methods for GATs adjust the attention coefficient of neighbors by improving the attention mechanism [47, 48], which can be considered as a deduction of GNN Architecture Tactic.
@@
In general, existing methods primarily focus on addressing a singular issue: how to enhance the aggregation of information from other nodes? The GNN Architecture Tactic targets at the Aggregation Mechanism, while the Graph Structure Tactic targets at the Aggregation Object. However, the Aggregation operation is derived from strong homophily hypothesis, which is not satisfied by heterophilic graphs. Therefore, modifying the Aggregation operation is not an essential way. Besides, more similar aggregation objects which can improve the homophilic ratio is not a necessary condition to explain the poor performance on heterophilic graphs of GNNs [24], and has the potential over-smoothing crisis.
@@
Contrary to the emphasis on Aggregation, we propose a new insight with the consideraion of the mechanism of GATs: enabling the central node to concentrate on itself while avoiding the distraction caused by a subset of neighbors can improve the discrimination ability of GATs on heterophilic graphs. We depict a toy example in Figure 1. For heterophilic graphs, the high proportion of inter-class edges lead to the updated representation of the central node deviating from the distribution of its own class, even when similar neighbors are assigned higher weights; Whereas after graph trimming, despite the increased homophilic ratio, the representation of the central node deviates less and is classified correctly due to higher self-attention and lower distraction. Since the removed nodes contribute to the decreased self-attention of the central node, and removing them helps prevent the occurring deviation, we refer to these nodes as Distraction Neighbors. They mathematically equal to the neighbors with High DE on the central node.
@@
Figure 1: Toy example: The comparison between low self-attention and high self-attention for the central node.
@@
In order to identify and remove Distraction Neighbors, we need to measure the DE of neighboring nodes on the central node, that is, the effect on the attention distribution of the central node generated from its neighbors. Therefore, two crucial questions need to be answered.
@@
Question 1: What is the basic unit of Distraction Neighbors when influencing the attention learning of the central node?
@@
Answer 1: Using two two heterophilic graphs as an example, we intervene the Local Neighbor Distribution (LND) of nodes and design three control groups to explore the effect of LND on the discrimination ability of the central node (Section 4.1). Experiments reveal that the semantic information provided by nodes of the same class is similar, referred to as Class-Level Semantic. Based on this observation, we introduce the concept of Class-Level Semantic Cluster and further propose the Class-Level Semantic Space hypothesis in (Section 4.2). In this hypothesis, we believe that neighbors of the same class have the similar impact on identifying the central node, therefore the basic unit for measuring DE should be class. It is more beneficial to get the genuine and stable effect of neighbors by treating the neighbors of the same class as a group.
@@
Module 1. Based on Answer 1, we design the Class-level Semantic Clustering Module, to pre-cluster local neighbors and get different Semantic Clusters for measuring their DE on the central node.
@@
Question 2: How much the Distraction Neighbors has impacted the attention learning of the central node?
@@
Answer 2: To better estimate the DE, we model DE as a type of causal effect. Specially, we formalize the influencing paths of neighboring nodes on the central node’s attention learning based on the working mechanism of GATs and construct a causal graph (Figure 2 and Figure 9). A causal graph is a graphical language capable of measuring causal effects through the manipulation of causal variables, referred to as intervention. Since the neighboring nodes influence the central node through two paths, we chose the Total Effect to estimate the overall causal effect.
@@
Module2. Based on Answer 2, we design the Total Effect Estimation Module, Intervene the LND of central node with Semantic Cluster as the basic unit, then calculate the TE from the changes of the central node’s attention distribution before and after intervention. Distraction Neighbors are identified and removed in accordance with TE and the corresponding trimmed graph are generated.
@@
    We propose a novel insight to enhance GAT’s discrimination ability on heterophilic graphs: maintaining the central node’s self-attention while avoiding distraction caused by neighbors. Instead of altering the architecture of GAT and searching for new neighbors globally, we use the attention distributed learned by GAT as signal to identify and remove Distraction Neighbors, which can be regarded as a trimming operation on the graph.
@@
    We propose a Causally graph Attention network for Trimming heterophilic graph (CAT) to improve the discrimination ability of GAT on heterophilic graphs. We employed three GATs as base model and conducted node classification experiments on seven datasets in three different sizes. Comparison Experiments show all base model can perform better within the CAT architecture. Ablation experiments and visualization experiments further validate the effectiveness.
@@
    We conduct pre-experiments and analyze the mechanism that how LND influence the attention learning of the central node based on our observations and background knowledge. We further formalize it into a causal graph where the causal effect of neighbors can be estimated with manipulations.
The rest of this paper is mainly organized as follows: in Section 2, we classify and summarize the existing methods on heterophilic graphs for both generous GNN and GAT-oriented, and list the differences between our methods and theirs. In Section 3, we introduce some important concepts and background knowledge needed for this paper, including the causal graphs derived from the background knowledge; In Section 4, we present the pre-experiments we did and the hypotheses we drew from the experiment results. We introduce our method in Section 5 and describe the dataset and experiments in Section 6. In Section 7 and Section 8, we discuss and conclude the work, fundamental issues that are inspiring and need to be answered in the future are also proposed.
@@ 2 Related Work
Strong homophily underlying graphs hypothesizes that connected nodes are similar, which is a necessity of GNNs. This principle is also widely acknowledged in various domains such as social networks [49] and citation networks [50, 51, 52]. Under this assumption, aggregating the information of neighbors gradually brings nodes of the same class closer in the representation space, thereby improving the discrimination ability of GNNs. However, when confronted with heterophilic graphs, the merit of GNNs may turn into a nightmare. The decline of GAT is especially obvious (Section 4.1). GNNs for heterophilic graphs have attracted increasing attention, and how to improve their discrimination ability poses significant challenges. A growing body of research has been developed to overcome the heterophilic challenges, and we categorized them into two groups based on their fundamental strategies: GNN Architecture Tactic and Graph Structure Tactic.
GNN Architecture Tactic. This kind of approach achieves better learning and fusion of neighbors’ information by altering the architecture of GNN.
@@
    Some methods aim to modify the aggregation operation in message passing. Ordered GNN [30] leverages a rooted-tree hierarchy aligning strategy to order the message passing, thereby achieving better fusion of information provided by nodes in different hops. NHGCN [29] designs a new metric, Neighborhood Homophily (NH) to group and aggregate the neighbors differently. LW-GCN [28] proposes a label-wise message passing mechanism that use pseudo labels to guide the aggregation of similar nodes and preserve the heterophilic contexts. DMP [26] takes the attributes as weak labels to measure attribute homophily rate, which can specify the attribute weight of edges for aggregation. CPGNN [27] incorporates an interpretable compatibility matrix for modeling the heterophily or homophily level, and uses the matrix to propagate and update the prior belief of each nodes. GGCN [31] proposes two strategies, structure-based and feature-based edge correction to adjust edges weights for aggregation.
@@
    Some methods aim to design different GNN layers and their relationships. Auto-HeG [33] builds a comprehensive GNN search space from which the optimal heterophilic GNN is selected. IIE-GNN [35] designs a GNN framework that contains seven blocks in four layers to enrich the intra-class information extraction. H2GCN [34] uses a combination of intermediate layers to concatenate the node representations from all previous layers, so as to capture local and global information better. GPR-GNN [32] combins a Generalized PageRank with GNN to learn weights of GNN layers for combination of intermediate layer representation.
@@
    Some methods aim to train GNNs in a new learning paradigm. HLCL [36] uses graph filters to generate augmented graph views and contrast the high-pass filter representation with the low-pass part for graph contrastive learning under heterophily. MVGE [37] builds two augmentation views with the input of ego features and aggregated features respectively, and forces the model to learn different graph signal through graph reconstruction task. SimP-GCN [38] employes self-supervised learning to capture the complex feature similarity and dissimilarity relations between nodes, which can help conduct the node similarity preserving aggregation.
Graph Structure Tactic. This kind of approach primarily involves modifying the connectivity of the graph, restructuring the graph by searching for more similar neighbors, and aggregating their information.
@@
    Some methods seek for similar neighbors from high-order neighbors. U-GCN [40] uses a multi-type convolution mechanism to capture and fuse the information from 1-hop, 2-hop and kNN neighbors. GPNN [39] adds the most relevant nodes from a large amount of multi-hop neighborhoods, and filters out irrelevant or noisy nodes from local neighborhoods.
@@
    Some methods search for nearest neighbors in a learned feature space. GEOM-GCN [42] learns a latent space and aggregate information from neighbors in latent space to strengthen the GCN’s ability to capture long-range dependencies in heterophilic graphs. Non-local GNNs [44] leverages attention mechanism to sort and find distant but informative nodes for conducting non-local aggregation. HOG-GCN [45] designs a novel propagation mechanism guided by the homophily degree between node pairs learned in the homophily degree matrix estimation module, to explore more similar node pairs. GCN-SL [41] uses spectral clustering to construct a re-connected graph according to the similarities between nodes and perform aggregation on the re-connected graph. A graph restructuring method [46] based on adaptive spectral clustering improves the node classification accuracy of GNN by improving graph homophily. DHGR [43] rewires graph by adding homophilic edges and pruning heterophilic edges, and the similarity of label/feature-distribution of node neighbors is adopted to determine the rewiring.
There are also researches focused on solutions for GATs, we called them GAT-oriented methods. HA-GAT [48] proposes a heterophily-aware attention scheme to assign weights for edges adaptively, and learns the local attention pattern of the central node by learning importance of each heterophilic edge type to get the final attention coefficient. GATv3 [47] proposes a new attention architecture to compute the attention coefficients between query and key, which improves the quality of the representations of nodes connected by heterophilic edges by introducing the representations learned by other GNNs. This type of approaches belongs to GNN Architecture Tactic.
Difference to GNNs for heterophilic graph. Our approach differs from the above GNNs for heterophilic graph in that it does not require altering the original GAT model or searching for new neighbors globally in the graph, but instead removes Distraction Neighbors in a graph trimming manner. It enables GAT to fully utilize its attention mechanism to improve its discrimination ability on heterophilic graphs.
Difference to GAT-oriented methods. Our approach is different from the above approaches targeted at GAT in that we do not need to change the attention mechanism of the original GAT model; instead, we make full use of the learned attention distribution of the original GAT model as a signal to find a better attention distribution. Therefore, our method is plug-and-play and applicable to any LAMP-driven GATs.

### v3

@@ Abstract
The local attention-guided message passing mechanism (LAMP) adopted in graph attention networks (GATs) can adaptively learn the importance of neighboring nodes and perform local aggregation better, thus demonstrating a stronger discrimination ability. However, existing GATs suffer from significant discrimination ability degradations in heterophilic graphs. The reason is that a high proportion of dissimilar neighbors can weaken the self-attention of the central node, resulting in the central node deviating from its similar nodes in the representation space. This type of influence caused by neighboring nodes is referred to as Distraction Effect (DE) in this paper. To estimate and weaken the DE induced by neighboring nodes, we propose a Causal graph Attention network for Trimming heterophilic graphs (CAT). To estimate the DE, since DE is generated through two paths, we adopt the total effect as the metric for estimating DE; To weaken the DE, we identify the neighbors with the highest DE (we call them Distraction Neighbors) and remove them. We adopt three representative GATs as the base model within the proposed CAT framework and conduct experiments on seven heterophilic datasets of three different sizes. Comparative experiments show that CAT can improve the node classification accuracies of all base GAT models. Ablation experiments and visualization further validate the enhanced discrimination ability of CATs. In addition, CAT is a plug-and-play framework and can be introduced to any LAMP-driven GAT because it learns a trimmed graph in the attention-learning stage, instead of modifying the model architecture or globally searching for new neighbors. The source code is available at https://github.com/GeoX-Lab/CAT.
@@ 1 Introduction
Graph neural networks (GNNs) are the most reliable and prevailing benchmark models for graph learning. With their effectiveness at representing irregular graph data, GNNs achieve state-of-the-art performance in tasks such as node classification, link prediction, graph classification, graph generation, and graph similarity calculation. They have also been widely applied in various fields such as recommendation systems, computer vision, natural language processing, molecular, and transportation. Their graph representation capability primarily stems from the ability to aggregate information [1], essentially following the message passing mechanism, which can build invariant input representations for the central node based on its neighbors. Existing GNNs utilize various aggregation operations following their fundamental assumptions about the influence of neighbors. However, they are all founded on the strong homophily hypothesis, obeying the rule that neighbors tend to be similar [2]. Among these GNNs, the graph attention network (GAT) [3] is a representative network that adaptively learns the importance of neighbors for aggregation through the local attention-guided message passing Mechanism (LAMP); therefore, it has the potential to achieve better performance on high-homophily graphs. However, the reverse of this situation is that the GATs’ performance decreases when addressing low-homophily graphs because assigning different aggregation weights under the smoothing principle leads to the failure to aggregate beneficial information and disrupt the raw features [4]. Experiments have shown that GNNs exhibit significant declines in node classification tasks [2] when the input graph is heterophilic, and we find that LAMP-driven GATs exhibit the most notable declines (as shown in Section 4.1). The primary reason for this phenomenon is the high proportion of dissimilar neighbors. Dissimilar neighbors influence the representation of the central node through their assigned attention levels and the weakened self-attention of the central node; both situations can result in the central node deviating from its similar nodes in the representation space. We refer to this impact of neighboring nodes on the central node as Distraction Effect (DE), which is generated through two paths (capturing the attention assigned to neighbors and reducing the self-attention of the central node). Improving the discrimination capabilities of GATs on heterophilic graphs poses a significant challenge.
@@
Great efforts have been devoted to improving the discrimination ability of general GNNs on heterophilic graphs, with only a few offering specific solutions for GATs. Based on the fundamental strategy targeted at GNN models or input graph data, these approaches can be categorized into two groups: GNN architecture-based tactics and graph structure-based tactics. GNN architecture-based tactics methods focus on modifying the GNN architecture to better utilize the information from neighboring nodes for aggregation. gNovel aggregation mechanisms have been proposed to adjust the weights of neighbors [5, 6, 7, 8, 9, 10, 4], and some works have fused information derived from different GNN layers in a new way [11, 12, 13, 14, 15, 16]. Self-supervised learning has also been adopted to capture more information from neighbors [17, 18, 19, 20, 21, 22, 23, 24, 25, 17]. Some methods focus on improving attention mechanisms [26, 27, 28]. Graph structure-based tactics methods, on the other way, focus on making heterophilic graphs more homophilic by obtaining more similar nodes for aggregation. These methods search for high-order neighbors [29, 30, 31, 32] or nearer neighbors in latent spaces [33, 34, 35, 36, 37, 38], forcing the central node closer to similar nodes in the representation space.
@@
In general, existing methods primarily focus on addressing a single issue: how to enhance the process of aggregating information from other nodes? GNN architecture-based tactics target at the aggregation mechanism, while graph structure-based tactics target at the aggregation source. The former involves the weights assignment, feature transformation, and learning paradigms for aggregation; while the latter proposes strategies for selecting the set of nodes for aggregation. They represent two different perspectives on improving aggregation respectively. However, the aggregation operation is derived from the strong homophily hypothesis, which is not satisfied by heterophilic graphs. Therefore, modifying the aggregation operation is not essential. In addition, searching for aggregation sources with higher similarity is not necessary for explaining the poor performance of GNNs on heterophilic graphs [2] and has the potential to cause oversmoothing.
@@
Contrary to the emphasis on aggregation, we propose a new insight concerning the mechanism of GATs: enabling the central node to concentrate on itself and avoiding the distraction during the aggregation can improve the discrimination ability of GATs on heterophilic graphs. We illustrate a representative example in Figure 1. For heterophilic graphs, a high proportion of interclass edges leads to the updated representation of the central node deviating from the distribution of its class, even when similar neighbors are assigned higher weights. After graph trimming, despite the decreased homophilic ratio, the representation of the central node deviates less and is classified correctly due to the higher self-attention and lower distraction level. Since the removed nodes contribute to the decreased self-attention of the central node, and removing them helps prevent deviations, we refer to these nodes as Distraction Neighbors. They are mathematically equal to the neighbors with high DEs.
@@
Figure 1: Toy example: A comparison between scenarios with low self-attention and high self-attention for the central node.
@@
To identify and remove Distraction Neighbors, we need to measure the DE of neighboring nodes on the central node, that is, the effect of the neighbors on the attention distribution of the central node. Therefore, two crucial questions must be answered.
@@
Question 1: What is the basic unit of Distraction Neighbors when influencing the attention-learning of the central node?
@@
Answer 1: Using two heterophilic graphs as an example, we intervene in the local neighbor distribution (LND) of the nodes and obtain three control groups to explore the effect of the LND on the discrimination ability of the central node (Section 4.1). Experiments reveal that nodes belonging to the same class provide similar semantic information; this kind of information is referred to as the Class-Level Semantic. Based on this observation, we introduce the concept of Class-Level Semantic Cluster and further propose the Class-Level Semantic Space hypothesis in (Section 4.2). According to this hypothesis, we believe that neighbors belonging to the same class have similar impacts on identifying the central node; therefore, the basic unit for measuring DE should be the class. It is more beneficial to obtain genuine and stable effects of neighbors by treating the neighbors belonging to the same class as a group.
@@
Module 1. Based on Answer 1, we design a Class-level Semantic Clustering Module, to precluster local neighbors and obtain different Semantic Clusters for measuring their DE on the central node.
@@
Question 2: To what extent do the Distraction Neighbors influence the attention-learning of the central node?
@@
Answer 2: To better estimate the DE, we model the DE as a type of causal effect. Specifically, we formalize the influencing paths of neighboring nodes on the attention-learning process of the central node based on the working mechanism of the GAT and construct causal graphs (Figure 2 and Figure 9). Since the neighboring nodes influence the central node through two paths, we chose the total effect to estimate the overall causal effect.
@@
Module2. Based on Answer 2, we design a Total Effect Estimation Module, to intervene in the LND of central nodes with Semantic Cluster as the basic unit, and then calculate the TE from the changes in the attention distribution of the central node before and after the intervention. Distraction Neighbors are identified and removed according to the TE, and a corresponding trimmed graph is generated.
@@
    We propose a novel insight for enhancing the discrimination ability of GATs on heterophilic graphs: maintaining the self-attention of the central node and avoiding distraction caused by neighbors. Instead of altering the architecture of the GAT or searching for new neighbors globally, we use the attention distribution learned by GAT to identify and remove Distraction Neighbors, which can be regarded as performing a trimming operation on the graph.
@@
    We propose a Causal graph Attention network for Trimming heterophilic graphs (CAT), to improve the discrimination ability of GATs for heterophilic graphs. We employ three GATs as the base model and conduct node classification experiments on seven datasets of three sizes. Comparison experiments, ablation experiments, and visualization experiments validate the effectiveness of CAT.
@@
    We conduct pre-experiments and investigate the mechanism by which the LND influences the attention-learning of the central node based on our observations and background knowledge. We further formalize this idea into causal graphs.
The remainder of this paper is organized as follows: in Section 2, we classify and summarize existing GNN methods for heterophilic graphs. In Section 3, we introduce important concepts and background knowledge needed for this paper, including the causal graphs derived from the background knowledge; In Section 4, we present the pre-experiments and the hypotheses we drew from them. We introduce our method in Section 5 and describe the dataset and experiments in Section 6. In Section 7 and Section 8, we discuss and conclude this work, fundamental issues that need further investigation are also raised.
@@ 2 Related Work
The strong homophily assumption underlying graphs indicates that connected nodes are similar, which is a necessity of GNNs. This principle is also widely acknowledged in various domains such as social networks and citation networks. Under this assumption, aggregating the information of neighbors gradually brings nodes belonging to the same class closer in the representation space, thereby improving the discrimination ability of GNNs. However, when confronted with heterophilic graphs, the merits of GNNs may not be realized. The declines exhibited by GATs are particularly pronounced (Section 4.1). GNNs for heterophilic graphs have attracted increasing attention, and we categorize the approaches aimed at overcoming these challenges into two groups based on their fundamental strategies: GNN architecture-based tactics and graph structure-based tactics.
GNN architecture-based tactics. The fundamental question addressed by methods in this line is how to more effectively aggregate information from neighbors. Therefore, these methods design and build various GNN architectures to better learn and fuse the information of neighbors.
@@
    Some methods aim to modify the aggregation operation in message passing. Various kinds of graph information are leveraged to guide the neighbor propagation process, where aggregation weights are learned to enhance similar features and weaken dissimilar features. An ordered GNN [5] leverages a rooted-tree hierarchy aligning strategy to order message passing, thereby achieving better fusing of information provided by nodes in different hops. NHGCN [6] employs a new metric, Neighborhood Homophily (NH) to group and aggregate the neighbors differently. LW-GCN [7] proposes a labelwise message passing mechanism that uses pseudolabels to guide the aggregation of similar nodes and preserve heterophilic contexts. DMP [8] takes attributes as weak labels to measure the attribute homophily rate, and to specify the attribute weights of the edges for aggregation. CPGNN [9] incorporates an interpretable compatibility matrix for modelling the heterophily or homophily level, and uses this matrix to propagate and update the prior belief of each node. GGCN [10] proposes two strategies, structure-based and feature-based edge correction to adjust the edge weights for aggregation. SAGNN [4] implements a sign attention mechanism to adaptively learn the weights of neighbors, which aggregates positive and negative information for neighbors within the same class and in different classes, respectively.
@@
    Some methods aim to design different GNN layers and determine their relationships. Because different layers in a GNN can encode different levels of node features, specific information can be learned by combining different intermediate layers. Auto-HeG [11] builds a comprehensive GNN search space from which the optimal heterophilic GNN is selected. IIE-GNN [12] designs a GNN framework that contains seven blocks in four layers to enrich the intraclass information extraction process. H2GCN [13] uses a combination of intermediate layers to concatenate the node representations derived from all previous layers, thereby better capturing local and global information. GPR-GNN [14] combines a Generalized PageRank algorithm with a GNN to learn the weights of GNN layers for combination with the intermediate layer representation. PCNet [15] employs a PC-Conv to perform both homophilic and heterophilic aggregation of node information, and SPCNet [16] further improves this approach.
@@
    Some methods aim to train GNNs in new learning paradigms. Self-supervised learning is a new paradigm that can help models learn better representations by leveraging the intrinsic structure of graphs. Multiview learning is a popular method that learns from multiple views via contrastive learning or invariance regularization [17], to capture rich information from unlabelled nodes. HLCL [18] and PolyGCL [19] use graph filters to generate augmented graph views and contrast the high-pass filter representation with the low-pass representation for conducting graph contrastive learning under heterophily. MVGE [20] builds two augmentation views with input ego features and aggregated features, and forces the model to learn different graph signals through a graph reconstruction task. GREET [21] trains an edge discriminator to augment homophilic and heterophilic views, and then uses a dual-channel contrastive loss to learn node representations. LHS [22] adopts a self-expressive generator to induce a latent homophilic structure via multinode interactions and iteratively refines the latent structure with a dual-view contrastive learner. MUSE [23] performs cross-view feature fusion across semantic and contextual views and learns perturbation-invariant representations via contrastive learning. SimP-GCN [24] employs a contrastive pretext task to capture the complex similar and dissimilar feature relations between nodes, which can help the method conduct node similarity-preserving aggregation. A multiresolution graph contrastive learning method [25] has been proposed that learns resolution invariant representations from graph augmentations constructed by diffusion wavelet filters. HGRL [17] adopts four types of graph augmentations and two pretext tasks to capture graph properties.
  4. 4.
    Some methods focus on GAT solutions, which are referred to as GAT-oriented methods. These approaches consider the characteristics of the GAT and perform better aggregation by proposing novel attention mechanisms. HA-GAT [26] utilizes a heterophily-aware attention scheme to adaptively assign weights for edges, and learns the local attention pattern of the central node by learning the importance of each heterophilic edge type. GATv3 [27] implements a new attention architecture to compute the attention coefficients between the query and key, which optimizes the representations of nodes by introducing representations learned by other GNNs. DGAT [28] leverages the diffusion distance to detect noisy neighbors and rewires heterophilic graphs, and proposes global directional attention to capture long-range neighborhood information.
Graph structure-based tactics. The core question behind this type of approach is how to select the neighbors that can provide beneficial information for aggregation. Therefore, these methods primarily involve restructuring a meaningful graph to connect more similar neighbors and then aggregating their information.
@@
    Some methods seek similar neighbors from high-order neighbors. With the experimental observation [29] that high-order neighborhoods may have higher homophily ratios, aggregating information from higher-order neighborhoods can lead to satisfactory performance. U-GCN [29] uses a multitype convolution mechanism to capture and fuse the information from 1-hop, 2-hop, and kNN neighbors. GPNN [30] adds the most relevant nodes from a large number of multihop neighborhoods, and filters out irrelevant or noisy nodes from the local neighborhoods. PathMLP [31] designs a similarity-based path sampling strategy guided by hop-by-hop similarity to conduct homophilic path aggregation. SFA-HGNN [32] uses high-order random walks to select and aggregate distant nodes.
@@
    Some methods search for nearest neighbors in a learned feature space. Graph representation learning methods aim to embed graphs into a latent space that approximates the inherent distribution space of the input data as closely as possible. Therefore, neighbors in this latent space contribute to better central node representations. GEOM-GCN [33] learns a latent space and aggregates information from neighbors in the latent space to strengthen the ability of GCN to capture long-range dependencies in heterophilic graphs. Non-local GNN [34] leverages attention mechanisms to sort and find distant but informative nodes for conducting nonlocal aggregation. HOG-GCN [35] designs a novel propagation mechanism guided by the homophily degree between node pairs learned in the homophily degree matrix estimation module. GCN-SL [36] uses spectral clustering to construct a reconnected graph according to the similarities between nodes and performs aggregation on it. A graph restructuring method [37] based on adaptive spectral clustering improves the node classification accuracy of GNNs by improving graph homophily. DHGR [38] rewires the graph by adding homophilic edges and pruning heterophilic edges, and the similarity of label/feature distribution of node neighbors is adopted to determine the rewiring strategy.
Our approach differs from the above GNNs for heterophilic graphs in that it does not require alterations of the original GNN models or global searching for new neighbors, but instead removes Distraction Neighbors via graph trimming. We make full use of the attention distribution learned by the original GAT models as signals, to find a better attention distribution. Therefore, our method is plug-and-play and is applicable to any LAMP-driven GAT.